OpenStudy (anonymous):

Please Help ME! x^2-3x-54/x^2-18x+81multiplied by x^2+12x+36/x+6?

jhonyy9 (jhonyy9):

(x-9)(x+6)/(x-9)(x-9) * (x+6)(x+6)/(x+6)

jhonyy9 (jhonyy9):

x+6 ---- * (x+6) x-9

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

what does that mean?

OpenStudy (anonymous):

x+6 over x-9?

jhonyy9 (jhonyy9):

x2 +12x +36 / x-9 = x+21 x2 -9x ------- 0 21x +36 21x - 189 ----------- 0 225

jhonyy9 (jhonyy9):

(x+6) squared -------- (x-9)

OpenStudy (anonymous):

now im confused at what you put up there but thanks :)

OpenStudy (anonymous):

\[\frac{x^2-3x-54}{x^2-18x+81}=\frac{(x-9)(x+6)}{(x-9)(x-9)}=\frac{x+6}{x-9}\]

OpenStudy (anonymous):

is the x+6 squared or not?

OpenStudy (anonymous):

\[\frac{x^2+12x+36}{x+6}=\frac{(x+6)^2}{x+6}=x+6\]

OpenStudy (anonymous):

multiply them together get \[\frac{(x+6)(x+6)}{x-9}\]

OpenStudy (anonymous):

or if you prefer \[\frac{(x+6)^2}{x-9}\]